OpenStudy (yttrium):

How far from the earth must a body be along line toward the sun so that the gravitational pull of the sun balances that of the earth? Earth-to-sun distance is 9.3x10^7 mi; mass of the sun is 3.24x10^5 times the mass of earth.

OpenStudy (masumanwar):

The mass should be at a distance where the gravitational pull due to sun on that mass according to the inverse square law of gravitational force should be equals to that of earth if the mass is specified then and assuming the distance as x we may calculate that position in the earth to sun joining line

OpenStudy (yttrium):

Can you elaborate it?

OpenStudy (anonymous):

\[G \frac{ M _{Sun}m }{ x ^{2} } = G \frac{ M _{Earth} m}{ (9.3\times 10^{7} - x)^{2} }\] G and m will be cancelled. and solve for x.

OpenStudy (anonymous):

like @masumanwar said